I don't understand how to do this : 3|x+6| = 36
Answer by MathLover1(20849)


3|x+6| = 36
|x+6| = 36/3
|x+6| = 12
so, x+6 = 12 or x+6 = -12
if x+6 = 12...->...x = 12-6..->...x =6
if x+6 = -12...->...x = -12-6..->...x =-18